# CareBot: A Pioneering Full-Process Open-Source Medical Language Model

### v1

@@ Abstract
Recently, both closed-source LLMs and open-source communities have made significant strides, outperforming humans in various general domains. However, their performance in specific professional domains such as medicine, especially within the open-source community, remains suboptimal due to the complexity of medical knowledge. In this paper, we propose CareBot, a bilingual medical LLM, which leverages a comprehensive approach integrating continuous pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning with human feedback (RLHF). Our novel two-stage CPT method, comprising Stable CPT and Boost CPT, effectively bridges the gap between general and domain-specific data, facilitating a smooth transition from pre-training to fine-tuning and enhancing domain knowledge progressively. We also introduce DataRater, a model designed to assess data quality during CPT, ensuring that the training data is both accurate and relevant. For SFT, we develope a large and diverse bilingual dataset, along with ConFilter, a metric to enhance multi-turn dialogue quality, which is crucial to improving the model’s ability to handle more complex dialogues. The combination of high-quality data sources and innovative techniques significantly improves CareBot’s performance across a range of medical applications. Our rigorous evaluations on Chinese and English benchmarks confirm CareBot’s effectiveness in medical consultation and education. These advancements not only address current limitations in medical LLMs but also set a new standard for developing effective and reliable open-source models in the medical domain. We will open-source the datasets and models later, contributing valuable resources to the research community.
@@ Introduction
Currently, there are several medical LLMs available in this domain . However, most of these models rely solely on Supervised Fine-Tuning (SFT) (Zhang et al. 2023, 2024a; Han et al. 2023). As is well-known, pre-training is a critical phase for learning domain-specific knowledge, and depending exclusively on SFT results in the models that can only produce answers in a fixed format. Another approach attempts to integrate pre-training with supervised fine-tuning by converting pre-training data in specific domains into a unified format similar to SFT data, such as (instruction, output) pairs using GPT-3.5 (Chen et al. 2023). This method of synthesizing large amounts of data can lead to the inclusion of significant amounts of incorrect knowledge that aligns with GPT-4 but diverges from human expertise, as well as high data synthesis costs. Furthermore, Yang et al. (2023b) introduce Zhongjing for Chinese medicine, which employs to implement the pipeline training from pre-training, SFT, to RLHF. However, this approach involves two phases of transformation for the base model, which may lead to issues such as catastrophic forgetting or model degradation (Cheng, Huang, and Wei 2024). Additionally, previous efforts have predominantly focused on data construction during the SFT stage (Li et al. 2023a; Zhang et al. 2024b, 2023), while neglecting the importance of data construction during the continue pre-training (CPT) stage. Yet, a well-designed CPT data strategy is also crucial for inserting medical expertise into the model.
@@
After extensive training and optimization, we successfully develope the CareBot. We rigorously evaluate the performance of our model using widely used Chinese and English benchmarks in the medical domain. The experimental results demonstrate that CareBot excels in both medical consultation and medical teaching, validating that our constructed datasets significantly enhance the model’s performance across multiple dimensions. The main contributions of this paper are as follows: (1) We design a novel two-stage CPT strategy that progressively and stably integrates domain knowledge into the LLM, and effectively addresses data bias and language imbalances in the original pre-training data. (2) We propose DataRater, a model for assessing data quality during CPT, ensuring that the data used for CPT is of high quality. (3) We construct a comprehensive open-source medical SFT dataset with high data diversity and data quality. Additionally, we develope ConFilter, a metric for measuring the correlation between multiple turns, to ensure the quality of multi-turn dialogues. (4) We conduct experiments across multiple Chinese and English benchmarks to validate the effectiveness and reliability of our training strategy and datasets.
construction process and each step is discussed sequentially to reflect the research workflow. The whole process is shown in Figure.
@@
Domain Classification Since the general pre-training corpus is sourced from diverse datasets and lacks clear domain labels, we first conduct domain classification to extract high-quality medical data. Specifically, we sample 40k data from the general corpus and use GPT-4 to perform two rounds of domain labeling to enhance accuracy. Data with inconsistent labels across the two rounds is removed, leaving us with 36k high-quality seed data. We observe that certain categories, such as artificial intelligence and computers, have long tails. To address this imbalance, we utilize GPT-4 to generate additional synthetic data for these long-tail categories. Finally, we design a domain classifier based on the bge-m3222https://huggingface.co/BAAI/bge-m3. We try a variety of multilingual pre-training models to train domain classifiers, details can be found in the Appendix A.
Rule-based Data Quality Filtering Quality filtering is a crucial step in processing the pre-training corpus. To eliminate noisy data, we use a rule-based filtering solution, including rules for removing data with insufficient tokens, excessive special characters, toxic content, and private information.
@@
To enhance the model’s ability to follow medical instructions and better adapt to specific medical scenarios, we conduct the supervised fine-tuning. This process involves using conversational-style data (comprising both queries and responses) to finetune the pretrained LLM. In the following sections, we will explore the details of data construction and training methods.
@@
Here, if $CF>1$, it means that historical information has a negative impact on current turn, that is, the correlation between contexts is very low. If $CF<1$, it means that historical information has a positive impact on current turn, that is, the correlation between contexts is high. However, too small $CF$ means that the context is highly repeated and the information is highly redundant. We also set a threshold to filter the data.
@@
In Figure 2, we evaluate our CPT model, CareBot, on seven common medical benchmarks. Considering that our goal is to train a medical model that performs well in both Chinese and English, we strive to improve the Chinese medical ability while ensuring that the English medical ability of the model is slightly reduced. We observe that for English benchmarks (MMLU-Med, PubMedQA, MedQA, MedMCQA), the performance of CareBot (Stable CPT) and CareBot (Stable CPT & Boost CPT) shows a slight decrease. This is expected, given that the LLaMA-8B-base model already has strong English capabilities. However, for Chinese benchmarks (C-Eval-Med, CMMLU-Med, CMB), our models demonstrate significant improvements, with particularly notable gains in models trained using the two-stage approach. This confirms that our two-stage CPT strategy effectively integrates medical domain knowledge into the model, resulting in robust enhancements to its Chinese medical capabilities.
@@
Multi-turn Dialogue In Table 2 and 3, we present the results for the multi-turn dialogue benchmarks CMtMedQA and CMB-clin, respectively. Overall, the performance of the two baselines is basically the same, with notable performance from HuatuoGPT and HuatuoGPT ll. Across the four dimensions, fluency and proficiency scores are particularly high, indicating coherent responses and a solid grasp of medical terminology by the medical LLMs. However, relevance and completeness scores are lower, suggesting room for improvement in providing highly relevant and comprehensive answers tailored to specific questions. Our model, CareBot-Chat, achieves strong performance across all dimensions, averaging scores of 4.35 and 4.22 respectively, with notable strengths in relevance and completeness. This underscores the effectiveness of our high-quality SFT dataset and our proposed multi-turn dialogue selection method, significantly enhancing the model’s contextual understanding and ensuring dialogue coherence and consistency. We further analyze the advantages of our models in multi-turn dialogues in Section Advantages of Multi-turn Dialogue.
Figure 3: The comparison of our CareBot-Chat’s predicted answers and other baselines’ predicted answers on single-turn dialogues from the Huatuo26M-test.
Figure 4: The comparison of our CareBot-Chat (RL)’s predicted answers and other baselines’ predicted answers on single-turn dialogues from the Huatuo26M-test.
Single-turn Dialogue Figure 3 displays the comparison results between our CareBot-Chat and various baselines on the single-turn dialogue benchmark Huatuo26M-test. It is evident that CareBot-Chat achieves comparable performance to ChatGPT and HuatuoGPT, indicating that the baseline HuatuoGPT performs similarly to ChatGPT under this evaluation framework. Moreover, our model significantly outperforms Zhongjing, MedAlpaca, and BioMistral. The lower performance of MedAlpaca and BioMistral can be attributed to their limited Chinese language capabilities and inadequate medical SFT data. It is noteworthy that Zhongjing, incorporating pre-training, SFT, and RLHF stages, performs poorly. This is likely due to the two training stages of the base model, leading to catastrophic forgetting or model degradation. Additionally, its instruction fine-tuning data is both limited and of inferior quality, which explains why a model with larger parameters than our CareBot performs substantially worse. However, we observe that our model only competes with HuatuoGPT ll-7B in the HuatuoEval framework. We will further discuss this phenomenon in Section Case Study. In Figure 4, we also compare our CareBot-Chat (RL) model with other baselines, and basically achieve the same performance as CareBot-Chat.
@@
To investigate the reasons behind our model’s strong performance in multi-turn dialogues, we analyze the quality of each turn of responses, as shown in Table 4 (Appendix G). Our analysis reveals that, except for the first turn, CareBot-Chat’s average score was 0.25 lower than that of HuatuoGPT II. However, in subsequent turns, CareBot-Chat consistently outperforms HuatuoGPT II. Furthermore, our advantage becomes increasingly pronounced as the number of dialogue turns increases. Specifically, in terms of fluency and proficiency, our model performs comparably to HuatuoGPT II. In contrast, HuatuoGPT II’s performance in relevance and completeness significantly deteriorates with more turns, while CareBot-Chat maintains stable performance. This highlights CareBot-Chat’s superior capability in multi-turn dialogues, demonstrating its better understanding of dialogue context. We attribute this advantage to our carefully curated multi-turn dialogue SFT data, which retains dialogues with contextual relevance without excessive repetition. Additionally, Figure 11 and 12 in Appendix G provide comparisons of responses from CareBot-Chat and HuatuoGPT II in the same multi-turn dialogue, further confirming our model’s effectiveness. Notably, in the first turn, HuatuoGPT II’s response is longer and more detailed than ours. However, from the second turn, while HuatuoGPT II’s answers remain detailed, they become increasingly irrelevant and lack coherence with the ongoing dialogue. Generally, our results strongly indicate that CareBot-Chat has a significant advantage in multi-turn dialogues, demonstrating superior contextual understanding and coherence.
@@
Table 13 (Appendix H) illustrates that our Carebot’s performance is 41.5% lower than HuatuoGPT II, indicating that our model does not hold a significant advantage over HuatuoGPT II. In this section, we present a case study. As depicted in Table 4, we provide an example where the output of CareBot-Chat is deemed inferior to that of HuatuoGPT II under the HuatuoEval evaluation framework. However, upon closer examination, we found there are some issues in HuatuoGPT II’s response, particularly in the statement ”such as 100% skim milk or low-fat milk. These milks usually contain more protein and calcium, and have lower sugar and fat content.” Firstly, this statement itself is an hallucination, that is, skim milk or low-fat milk does not contain more protein and calcium, but only has lower fat content. Furthermore, medical professionals confirm that whole milk is more suitable for infants due to its nutritional benefits, including fats crucial for development. This aligns with responses generated by GPT-4. Additionally, HuatuoGPT II often includes repetitive content such as ”It is better to choose milk designed specifically for children,” which, despite mimicking a doctor’s tone and offering longer responses, sometimes lacks relevance and completeness. This approach occasionally introduces ambiguities.

### v2

@@ Abstract
Recently, both closed-source and open-source LLMs have made significant strides, outperforming humans in various general domains. However, their performance in specific professional domains such as medicine, especially within the open-source community, remains suboptimal due to the complexity of medical knowledge. In this paper, we propose CareBot, a bilingual medical LLM, which leverages a comprehensive approach integrating continuous pre-training (CPT), supervised fine-tuning (SFT), and reinforcement learning with human feedback (RLHF). Our novel two-stage CPT method, comprising Stable CPT and Boost CPT, effectively bridges the gap between general and domain-specific data, facilitating a smooth transition from pre-training to fine-tuning and enhancing domain knowledge progressively. We also introduce DataRater, a model designed to assess data quality during CPT, ensuring that the training data is both accurate and relevant. For SFT, we develope a large and diverse bilingual dataset, along with ConFilter, a metric to enhance multi-turn dialogue quality, which is crucial to improving the model’s ability to handle more complex dialogues. The combination of high-quality data sources and innovative techniques significantly improves CareBot’s performance across a range of medical applications. Our rigorous evaluations on Chinese and English benchmarks confirm CareBot’s effectiveness in medical consultation and education. These advancements not only address current limitations in medical LLMs but also set a new standard for developing effective and reliable open-source models in the medical domain [^1] https://github.com/FlagOpen/CareBot.
@@ Introduction
Currently, there are several medical LLMs available in this domain . However, most of these models rely solely SFT (Zhang et al. 2023, 2024a; Han et al. 2023). As is well-known, pre-training is a critical phase for learning domain-specific knowledge, and depending exclusively on SFT results in models that can only produce answers in a fixed format. Another approach attempts to integrate pre-training with SFT by converting pre-training data in specific domains into a unified format similar to SFT data, such as (instruction, output) pairs using GPT-3.5 (Chen et al. 2023). This method of synthesizing large amounts of data can lead to the inclusion of significant amounts of incorrect knowledge that aligns with GPT-4 but diverges from human expertise, as well as high data synthesis costs. Furthermore, Yang et al. (2023b) introduce Zhongjing for Chinese medicine, which employs to implement the pipeline training from pre-training, SFT, to RLHF. However, this approach involves two phases of transformation for the base model, which may lead to issues such as catastrophic forgetting or model degradation (Cheng, Huang, and Wei 2024). Additionally, previous efforts have predominantly focused on data construction during the SFT stage (Li et al. 2023a; Zhang et al. 2024b), while neglecting the importance of data construction during the CPT stage. Yet, a well-designed CPT data strategy is also crucial for inserting medical expertise into the model.
@@
After extensive training and optimization, we successfully develope the CareBot. We rigorously evaluate the performance of our model using widely used Chinese and English benchmarks in medical domain. The experimental results demonstrate that CareBot excels in both medical consultation and teaching, validating that our constructed datasets significantly enhance the model’s performance across multiple dimensions. The main contributions of this paper are as follows: (1) We design a novel two-stage CPT strategy that progressively and stably integrates domain knowledge into the LLM, and effectively addresses data bias and language imbalances in the original pre-training data. (2) We propose DataRater, a model for assessing data quality during CPT, ensuring that the data used for CPT is of high quality. (3) We construct a comprehensive open-source medical SFT dataset with high data diversity and data quality. Additionally, we develope ConFilter, a metric for measuring the correlation between multiple turns, to ensure the quality of multi-turn dialogues. (4) We conduct experiments across multiple Chinese and English benchmarks to validate the effectiveness and reliability of our training strategy and datasets.
@@
Domain Classification Since the general pre-training corpus is sourced from diverse datasets and lacks clear domain labels, we first conduct domain classification to extract high-quality medical data. Specifically, we sample 40k data from general corpus and use GPT-4 to perform two rounds of domain labeling to enhance accuracy. Data with inconsistent labels across two rounds is removed, leaving us with 36k high-quality seed data. We observe that certain categories, such as artificial intelligence and computers, have long tails. To address this imbalance, we utilize GPT-4 to generate additional synthetic data for these long-tail categories. Finally, we design a domain classifier based on the bge-m3333https://huggingface.co/BAAI/bge-m3. We try a variety of multilingual pre-training models to train domain classifiers, details can be found in the Appendix A.
Rule-based Data Quality Filtering Quality filtering is a crucial step in processing pre-training corpus. To eliminate noisy data, we use a rule-based filtering solution, including rules for removing data with insufficient tokens, excessive special characters, toxic content, and private information.
@@
To enhance model’s ability to follow medical instructions and better adapt to specific medical scenarios, we conduct the SFT. This process involves using conversational-style data (comprising both queries and responses) to finetune the pretrained LLM. In the following sections, we will explore the details of data construction and training methods.
@@
Here, if $CF>1$, it means historical information has a negative impact on current turn, that is, the correlation between contexts is very low. If $CF<1$, it means historical information has a positive impact on current turn, that is, the correlation between contexts is high. However, too small $CF$ means that the context is highly repeated and the information is highly redundant. We also set a threshold to filter the data.
@@
In Figure 2, we evaluate our CPT model, CareBot, on seven common medical benchmarks. Considering that our goal is to train a medical model that performs well in both Chinese and English, we strive to improve Chinese medical ability while ensuring that English medical ability of the model is slightly reduced. We observe that for English benchmarks (MMLU-Med, PubMedQA, MedQA, MedMCQA), the performance of CareBot (Stable CPT) and CareBot (Stable CPT & Boost CPT) shows a slight decrease. This is expected, given that the LLaMA-8B-base model already has strong English capabilities. However, for Chinese benchmarks (C-Eval-Med, CMMLU-Med, CMB), our models demonstrate significant improvements, with particularly notable gains in models trained using the two-stage approach. This confirms that our two-stage CPT strategy effectively integrates medical domain knowledge into the model, resulting in robust enhancements to its Chinese medical capabilities.
@@
Multi-turn Dialogue In Table 2 and 3, we present the results for the multi-turn dialogue benchmarks CMtMedQA and CMB-clin, respectively. Overall, the performance of two baselines is basically the same, with notable performance from HuatuoGPT and HuatuoGPT ll. Across four dimensions, fluency and proficiency scores are particularly high, indicating coherent responses and a solid grasp of medical terminology by the medical LLMs. However, relevance and completeness scores are lower, suggesting room for improvement in providing highly relevant and comprehensive answers tailored to specific questions. Our model, CareBot-Chat, achieves strong performance across all dimensions, averaging scores of 4.35 and 4.22 respectively, with notable strengths in relevance and completeness. This underscores the effectiveness of our high-quality SFT dataset and our proposed multi-turn dialogue selection method, significantly enhancing the model’s contextual understanding and ensuring dialogue coherence and consistency. We further analyze the advantages of our models in multi-turn dialogues in Section Advantages of Multi-turn Dialogue.
(a) CareBot-Chat
(b) CareBot-Chat (RL)
Figure 3: Comparison of CareBot-Chat and CareBot-Chat (RL)’s predicted answers and other baselines’ predicted answers on single-turn dialogues from Huatuo26M-test.
Single-turn Dialogue Figure 3 (a) displays the comparison results between our CareBot-Chat and various baselines on the single-turn dialogue benchmark Huatuo26M-test. It is evident that CareBot-Chat achieves comparable performance to ChatGPT and HuatuoGPT, indicating that the baseline HuatuoGPT performs similarly to ChatGPT under this evaluation framework. Moreover, our model significantly outperforms Zhongjing, MedAlpaca, and BioMistral. The lower performance of MedAlpaca and BioMistral can be attributed to their limited Chinese language capabilities and inadequate medical SFT data. It is noteworthy that Zhongjing, incorporating pre-training, SFT, and RLHF stages, performs poorly. This is likely due to the two training stages of the base model, leading to catastrophic forgetting or model degradation. Additionally, its instruction fine-tuning data is both limited and of inferior quality, which explains why a model with larger parameters than our CareBot performs substantially worse. However, we observe that our model only competes with HuatuoGPT ll-7B in the HuatuoEval framework. We will further discuss this phenomenon in Section Case Study. In Figure 3 (b), we also compare our CareBot-Chat (RL) model with other baselines, and basically achieve the same performance as CareBot-Chat.
@@
To investigate the reasons behind our model’s strong performance in multi-turn dialogues, we analyze the quality of each turn of responses, as shown in Table 4 (Appendix G). Our analysis reveals that, except for the first turn, CareBot-Chat’s average score was 0.25 lower than that of HuatuoGPT II. However, in subsequent turns, CareBot-Chat consistently outperforms HuatuoGPT II. Furthermore, our advantage becomes increasingly pronounced as the number of dialogue turns increases. Specifically, in terms of fluency and proficiency, our model performs comparably to HuatuoGPT II. In contrast, HuatuoGPT II’s performance in relevance and completeness significantly deteriorates with more turns, while CareBot-Chat maintains stable performance. This highlights CareBot-Chat’s superior capability in multi-turn dialogues, demonstrating its better understanding of dialogue context. We attribute this advantage to our carefully curated multi-turn dialogue SFT data, which retains dialogues with contextual relevance without excessive repetition. Additionally, Figure 10 and 11 in Appendix G provide comparisons of responses from CareBot-Chat and HuatuoGPT II in the same multi-turn dialogue, further confirming our model’s effectiveness. Notably, in the first turn, HuatuoGPT II’s response is longer and more detailed than ours. However, from the second turn, while HuatuoGPT II’s answers remain detailed, they become increasingly irrelevant and lack coherence with the ongoing dialogue. Generally, our results strongly indicate that CareBot-Chat has a significant advantage in multi-turn dialogues, demonstrating superior contextual understanding and coherence.
@@
Figure 3 (a) illustrates our Carebot does not hold a significant advantage over HuatuoGPT II. Here, we present a case study. As depicted in Figure 12 (Appendix H), we provide an example where the output of CareBot-Chat is deemed inferior to that of HuatuoGPT II under the HuatuoEval evaluation framework. However, upon closer examination, we found there are some issues in HuatuoGPT II’s response, particularly in the statement ”such as 100% skim milk or low-fat milk. These milks usually contain more protein and calcium, and have lower sugar and fat content.” Firstly, this statement itself is an hallucination, that is, skim milk or low-fat milk does not contain more protein and calcium, but only has lower fat content. Furthermore, medical professionals confirm that whole milk is more suitable for infants due to its nutritional benefits, including fats crucial for development. This aligns with responses generated by GPT-4. Additionally, HuatuoGPT II often includes repetitive content such as ”It is better to choose milk designed specifically for children,” which, despite mimicking a doctor’s tone and offering longer responses, sometimes lacks relevance and completeness. This approach occasionally introduces ambiguities.
@@ Conclusion
Acknowledgments
We thank all anonymous reviewers. This work was supported by National Science and Technology Major Project No.2022ZD0116314.